OpenStudy (leeanm):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Write the equation of the line perpendicular to 3x +y = -8 that passes through (-3,1) . Write your answer in slope-intercept form. Show your work.

OpenStudy (triciaal):

suggest first write the equation in slope intercept form identify the slope for perpendicular line the slope is -1/slope of given line rephrased product of the slopes = -1 you now have the slope of the perpendicular and you know it passes through (-3,1) the slope is the same for any 2 points on that line

OpenStudy (leeanm):

@triciaal i dont understand

OpenStudy (leeanm):

@Directrix

OpenStudy (leeanm):

@BlazeRyder

OpenStudy (leeanm):

@563blackghost

OpenStudy (leeanm):

@mathmale

Directrix (directrix):

Follow @triciaal 's instructions 1) suggest first write the equation in slope intercept form 3x +y = -8 in slope-intercept form: y = -3x -8 2) identify the slope slope = -3 3) rephrased product of the slopes = -1 Slope of perpendicular = 1/3 because -3 * (1/3) = -1

Directrix (directrix):

You now have the slope of the perpendicular and you know it passes through (-3,1) What is the equation of the line with slope 1/3 that passes through the point ( -3, 1) Simplify this: y - 1 = (1/3) * (x - (-3) ) @leeanm

OpenStudy (leeanm):

@Directrix i dont know how to simplify that

OpenStudy (leeanm):

@likeabossssssss